OpenStudy (anonymous):

A girl moves down the hall at 1.2 m/s When she sees Peter coming, he begins to run is fear. After 3.2s he is moving 3.6m/s what is his acceleration?

OpenStudy (anonymous):

u=1.2m/s v=3.6 m/s t=3.2 s v=u+at plug the values and find the value of a

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

I do not get this

OpenStudy (anonymous):

3.6=1.2+3.2a 3.6-1.2=3.2a 2.4/3.2=a or a=2/3m/s^2

OpenStudy (anonymous):

so what would the answer be

OpenStudy (anonymous):

\[acceleration~is \frac{ 2 }{3 }m/s ^{2}\]

OpenStudy (anonymous):

I do not get this how you found out this

OpenStudy (anonymous):

correction \[It~is \frac{ 3 }{ 4 }m/s ^{2}\]

OpenStudy (anonymous):

How did you find out this

OpenStudy (anonymous):

could you draw it out

OpenStudy (anonymous):

it is a general formula in physics.

OpenStudy (anonymous):

what is the formula

OpenStudy (anonymous):

rate of change of velocity is called acceleration. Let initial velocity =u time=t velocity after time t=v change of velocity=v-u \[Rate~of~change~of~velocity=\frac{ v-u }{t }\] rate of change of velocity is called acceleration. \[Hence~a=\frac{ v-u }{ t },~or~v-u=at,~or~v=u+at\]

OpenStudy (anonymous):

oh ok thank you so much you saved my life :)

OpenStudy (anonymous):

np